A constant force of 245 N is applied at a 48.2° angle to a mass of 62.1 kg as shown below. If the mass moves at a constant speed of 3.28 ms–1, determine its coefficient of dynamic friction.  

Ans:0.206


1
Expert's answer
2022-04-24T17:19:16-0400

"N=mg+F\sin48.2°=792~N,"

"\mu=\frac{F\cos 48 .2°}N=0 .2."